# The Belle II Collaboration

### v1

@@ VII Conclusions
where the uncertainties are statistical, systematic and theoretical, respectively. This value is compatible within uncertainties with the average value obtained from measurements using $B\rightarrow\pi\ell\nu$ decays but it exceeds the HFLAV exclusive average [5]. The values of $|V_{ub}|$ extracted from the three phase-space regions are compared to the latest inclusive and exclusive HFLAV averages in Figure 9.
@@ Appendix A $B\rightarrow X_{c}\ell\nu$ suppression classifier input features
The shapes of the nine variables used to train the $B\rightarrow X_{c}\ell\nu$ suppression classifier (see Section III.5) are compared for $B\rightarrow X_{u}\ell\nu$ and $B\rightarrow X_{c}\ell\nu$ events in Figure 10. The performance of the classifier appears to be mostly driven by the event missing mass, the quality of ROE vertex fit and the missing mass in $D^{*}\to D\pi^{0}$ decays.

### v2

@@ VII Conclusions
where the uncertainties are statistical, systematic and theoretical, respectively. The GGOU framework is used to allow a consistent comparison with the inclusive $|V_{ub}|$ average reported by HFLAV, with which our result is compatible [13]. The measured value also agrees within uncertainties with the average value obtained from measurements using $B\rightarrow\pi\ell\nu$ decays but it exceeds the HFLAV exclusive average. The values of $|V_{ub}|$ extracted from the three phase-space regions are compared to the latest inclusive and exclusive HFLAV averages in Fig. 9.
@@ Appendix A $B\rightarrow X_{c}\ell\nu$ suppression classifier input features
The shapes of the nine variables used to train the $B\rightarrow X_{c}\ell\nu$ suppression classifier (see Sec. III.5) are compared for $B\rightarrow X_{u}\ell\nu$ and $B\rightarrow X_{c}\ell\nu$ events in Fig. 10. The performance of the classifier appears to be mostly driven by the event missing mass, the quality of ROE vertex fit and the missing mass in $D^{*}\to D\pi^{0}$ decays.